# Co-Engel graphs of certain finite non-Engel groups

### v3

@@ 4. Realization of $\mathop{\mathcal{E}_{c}^{-}}(G)$
From this point on, we turn our attention to properties of co-Engel graphs. In this section we realize certain graphs as $\mathop{\mathcal{E}_{c}^{-}}(G)$ for certain finite groups. Let $K_{n}$ be the complete graph on $n$ vertices and $K_{n_{1},n_{2},\dots,n_{m}}$ be the complete $m$-partite graph with parts of size $n_{1},n_{2},\dots,n_{m}$. We write $K_{m\cdot n}=K_{n_{1},n_{2},\dots,n_{m}}$ if $n_{1}=n_{2}=\cdots=n_{m}=n$.
@@
It is not difficult to see that if $G$ denotes one of the groups $D_{2^{t+1}m}$ and $Q_{2^{t+1}m}$, then $L(G)=\mathop{\mathrm{Fit}}(G)=\langle y\rangle$, where $\mathop{\mathrm{Fit}}(G)$ is the Fitting subgroup of $G$ (see [3]). Thus $G\setminus L(G)=\{x,xy,\dots,xy^{2^{t}m-1}\}$. We have
@@
We also have $a^{i}b^{t}a^{j}b^{s}=a^{i+j}b^{s+tr^{j}}$. Therefore, if $i\neq 0$ then $(a^{i}b^{j})^{p}=1$. Note that the group $F_{p,q}$ is the union of its subgroups $H_{b}=\langle b\rangle=\mathop{\mathrm{Fit}}(F_{p,q})=L(F_{p,q})$ and
@@ 5. Genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$
In this section, we determine the genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$ for the groups considered in Section 4. It is well-known that
$K_{n}=\left\lceil\frac{(n-3)(n-4)}{12}\right\rceil$ (1)
@@
$K_{m,n}=\left\lceil\frac{(m-2)(n-2)}{4}\right\rceil.$ (2)
@@
Proof.
By Theorem 4.1 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$, where $G=D_{2^{t+1}m}$ and $Q_{2^{t+1}m}$. Therefore, putting $a=m$ and $b=2^{t}$ in Theorem 5.1, we get
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{m(m-1)}{2}\left\lceil\frac{(2^{t}-2)^{2}}{4}\right\rceil+\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil.$
Hence, the first part of the result follows. The second part follows from (1) using the fact that $\gamma(\mathop{\mathcal{E}_{c}^{-}}(D_{2m}))=\gamma(K_{m})$. ∎
@@
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(F_{p,q})\cong K_{q\cdot(p-1)}$. If $p=2$ then $\mathop{\mathcal{E}_{c}^{-}}(F_{p,q})\cong K_{q}$ noting that $K_{q\cdot 1}=K_{q}$. Therefore, the result follows from (1).
If $p\geq 3$ then, putting $a=q$ and b=p1b=p-1 in Theorem 5.1, we get
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(F_{p,q}))=\frac{q(q-1)}{2}\left\lceil\frac{((p-1)-2)^{2}}{4}\right\rceil+\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil.$
Hence, the result follows. ∎
@@
Proof.
By Theorem 4.3 we have $\mathop{\mathcal{E}_{c}^{-}}(X)\cong K_{lm\cdot n}$. Hence, the result follows from Theorem 5.1, putting $a=lm$ and $b=n$. ∎
@@
Proof.
Note that $K_{m}\cong K_{m\cdot 1}$. By Theorem 4.3, we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{lm\cdot 1}\cong K_{lm}$. Hence, the result follows from (1). ∎
@@
Proof.
By Theorem 4.1 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$ if $G=D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$, and $\mathop{\mathcal{E}_{c}^{-}}(D_{2m})\cong K_{m}$. In view of this, it is easy to see that
If $|H|=l$ and $K=D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$ then, by Theorem 4.3, $\mathop{\mathcal{E}_{c}^{-}}(H\times K)\cong K_{(}lm\cdot 2^{t})$ and $\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2m})\cong K_{lm}$. By Theorem 5.1 we have
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times K))=\frac{lm(lm-1)}{2}\left\lceil\frac{(2^{t}-2)^{2}}{4}\right\rceil+\left\lceil\frac{(lm-3)(lm-4)}{12}\right\rceil.$
Hence we get the required expression for $\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times K))$. The expression for $\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2m}))$ can be obtained from (1). ∎
@@
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(F_{p,q})=K_{q\cdot(p-1)}$. In view of this and Theorem 4.3, it follows that $\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q})\cong K_{lq\cdot(p-1)}$.
Putting $a=lq$ and b=p1b=p-1 in Theorem 5.1, we have
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q}))=\frac{lq(lq-1)}{2}\left\lceil\frac{((p-1)-2)^{2}}{4}\right\rceil+\left\lceil\frac{(lq-3)(lq-4)}{12}\right\rceil.$
Hence, we get the required expression for $\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q}))$. ∎
@@
Since $\omega(\mathop{\mathcal{E}_{c}^{-}}(G))\leq 4$, by [1, Proposition 1.4] and [2, Theorem 1.2], we have $\frac{G}{Z^{*}(G)}\cong D_{6}$ or $A_{4}$.
Suppose $\frac{G}{Z^{*}(G)}\cong D_{6}$. Let $\bar{G}=\frac{G}{Z^{*}(G)}$ and $\bar{x}=xZ^{*}(G)$ for every $x\in G$. Since $\bar{G}\cong D_{6}$, we have $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\cong\mathop{\mathcal{E}_{c}^{-}}(D_{6})$. Let $\phi:\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\longrightarrow\mathop{\mathcal{E}_{c}^{-}}(D_{6})$ be a graph isomorphism. Let $a,b\in D_{6}$ such that $D_{6}=\langle a,b\rangle$ and $a^{3}=b^{2}=1$. Then $\{b,ab,a^{2}b\}$ forms a clique in $\mathop{\mathcal{E}_{c}^{-}}(D_{6})$. Therefore $\{\bar{x}_{1}=\phi^{-1}(b),\bar{x}_{2}=\phi^{-1}(ab),\bar{x}_{3}=\phi^{-1}(a^{2}b)\}$ forms a clique of $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})$. By Theorem 3.2, every elements of $x_{i}Z^{*}(G)$ is adjacent to every elements of $x_{j}Z^{*}(G)$, for $i,j\in\{0,1,2\},i\neq j$. Thus $K_{m,m,m}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$, where $m=|Z^{*}(G)|$. Therefore, by (3), we have
@@
If $m=2$, then $G$ is a non-nilpotent group of order $12$ with $|Z^{*}(G)|=2$. Thus $G\cong D_{12},Q_{12}$. Therefore, $\mathcal{E}_{G}[G\setminus L(G)]$ is planar
If $m=3$, then $G$ is a non-nilpotent graph of order $18$ with $|Z^{*}(G)|=3$. Thus $G\cong C_{3}\times D_{6}$. Here $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})\cong K_{3,3,3}$ and $\gamma(K_{3,3,3})=\frac{(3-1)(3-2)}{2}=1$. Thus $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})$ is toroidal. An embedding of $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})$ on a torus is shown in Figure 1, where $C_{3}=\langle a\rangle$.
@@
Suppose $\frac{G}{Z^{*}(G)}\cong A_{4}$. Suppose that $|Z^{*}(G)|\geq 2$. Let $\bar{G}=\frac{G}{Z^{*}(G)}$ and $\bar{x}=xZ^{*}(G)$ for every $x\in G$. Since $\bar{G}\cong A_{4}$, we have $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\cong\mathop{\mathcal{E}_{c}^{-}}(A_{4})$. Let $\phi:\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\longrightarrow\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ be a graph isomorphism. As we see in Figure 2, $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ has a subgraph isomorphic to $K_{4,4}$ with parts $H=\{(2,3,4),(1,2,4),(2,4,3),(1,4,2)\}$ and $K=\{(1,2,3),(1,3,4),(1,3,2),(1,4,3)\}$. Therefore, $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})$ has a subgraph isomorphic to $K_{4,4}$ with parts
@@
Let $z\in Z^{*}(G),z\neq 1$. Then, by Theorem 3.2, every elements of $\bar{H}\cup z\bar{H}$ is adjacent to every elements of $\bar{K}\cup z\bar{K}$, showing that $K_{8,8}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$.
Thus $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq\gamma(K_{8,8})=9$, a contradiction. Thus $|Z^{*}(G)|=1$ and so $G\cong A_{4}$. An embedding of the graph $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ on a torus is shown in Figure 3. This completes the proof. ∎
@@
By [1, Proposition 1.4] and [2, Theorem 1.2], we have $G/Z^{*}(G)\cong D_{6}$ or $A_{4}$. Suppose $G/Z^{*}(G)\cong A_{4}$. As seen in Figure 2, $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ has a subgraph isomorphic to $K_{4,4}$. Thus $\mathop{\mathcal{E}_{c}^{-}}(G/Z^{*}(G))$ has a subgraph isomorphic to $K_{4,4}$. Note that $|v(\mathop{\mathcal{E}_{c}^{-}}(G/Z^{*}(G)))|=|v(\mathop{\mathcal{E}_{c}^{-}}(A_{4}))|=8$. Let $v(\mathop{\mathcal{E}_{c}^{-}}(G/Z^{*}(G)))=\{\bar{a_{1}},\bar{a_{2}},\dots,\bar{a_{8}}\}$. Then the induced subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$ by the set $\{a_{1},a_{2},\dots,a_{8}\}$ has a subgraph isomorphic to $K_{4,4}$. Thus $\bar{\gamma}(\mathop{\mathcal{E}_{c}^{-}}(G))\geq\bar{\gamma}(K_{4,4})=2$, a contradiction. Thus $G/Z^{*}(G)\cong D_{6}$.
Let $m=|Z^{*}(G)|$. As seen in the proof of Theorem 5.12, $K_{m,m,m}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$. Thus if $m\geq 3$, then $K_{6,3}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$. But $\bar{\gamma}(K_{6,3})=2$. Therefore, $m\leq 2$. If $m=1$, then $G\cong D_{6}$ and if $m=2$, then $G\cong D_{12},Q_{12}$. Note that $\mathop{\mathcal{E}_{c}^{-}}(D_{6})\cong K_{3}$ is projective. Also, from Theorem 4.1, $\mathop{\mathcal{E}_{c}^{-}}(D_{12})\cong\mathop{\mathcal{E}_{c}^{-}}(Q_{12})\cong K_{3\cdot 2}$ which is projective. This completes the proof. ∎
@@ 6. Various spectra and energies of $\mathop{\mathcal{E}_{c}^{-}}(G)$
If $m\geq 3$ is odd and $G\cong D_{2m}$ then, by Theorem 5.6, we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m}$. Therefore, we get the following result.
@@
Proof.
By Theorem 5.6 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$. Using Lemma 6.2 we get
$P(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x^{m(2^{t}-1)}(x+2^{t})^{m-1}\left(x-2^{t}(m-1)\right),$
$P_{L}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x(x-2^{t}(m-1))^{m(2^{t}-1)}(x-2^{t}m)^{m-1}\text{ and }$
$P_{Q}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=(x-2^{t}(m-1))^{m(2^{t}-1)}(x-2^{t}(m-2))^{m-1}(x-2^{t+1}(m-1)).$
Therefore, $\operatorname{Spec}(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{m(2^{t}-1)},(-2^{t})^{m-1},\left(2^{t}(m-1)\right)^{1}\right\}$ and so
$E(\mathop{\mathcal{E}_{c}^{-}}(G))=m(2^{t}-1)\times 0+(m-1)2^{t}+2^{t}(m-1)=2^{t+1}(m-1).$
Also, $\operatorname{L-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{1},(2^{t}(m-1))^{m(2^{t}-1)},(2^{t}m)^{m-1}\right\}$ and
$\operatorname{Q-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(2^{t}(m-1))^{m(2^{t}-1)},(2^{t}(m-2))^{m-1},(2^{t+1}(m-1))^{1}\right\}$.
We have ${|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{2^{t}m(2^{t}m-1)}{2}-\frac{2^{t}m(2^{t}-1)}{2}=\frac{2^{t}m(2^{t}m-2^{t})}{2}$ and so $\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}$ $2^{t}(m-1)$. Therefore,
$\left|0-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=\left|-2^{t}(m-1)\right|=2^{t}(m-1),$
$\left|2^{t}(m-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t}(m-1)-2^{t}(m-1)\right|=0,$
and
$\left|2^{t}m-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t}m-2^{t}(m-1)\right|=2^{t}.$
Therefore,
$LE(\mathop{\mathcal{E}_{c}^{-}}(G))$ $=1\times 2^{t}(m-1)+m(2^{t}-1)\times 0+(m-1)\times 2^{t}$
$=2^{t+1}(m-1).$
Also,
$\left|2^{t}(m-2)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t}(m-2)-2^{t}(m-1)\right|=2^{t},$
and
$\left|2^{t+1}(m-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t+1}(m-1)-2^{t}(m-1)\right|$
$=2^{t}(m-1).$
Therefore
$LE^{+}(\mathop{\mathcal{E}_{c}^{-}}(G))$ $=m(2^{t}-1)\times 0+(m-1)\times 2^{t}+1\times 2^{t}(m-1)$
$=2^{t+1}(m-1).$
@@
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{q\cdot(p-1)}$. Using Lemma 6.2 we get
$P(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x^{q(p-2)}(x+(p-1))^{q-1}\left(x-(p-1)(q-1)\right),$
$P_{L}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x(x-(p-1)(q-1))^{q(p-2)}(x-q(p-1))^{q-1}\text{ and }$
$P_{Q}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=(x-(p-1)(q-1))^{q(p-2)}(x-(p-1)(q-2))^{q-1}(x-2(p-1)(q-1)).$
Therefore, $\operatorname{Spec}(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{q(p-2)},(-(p-1))^{q-1},\left((p-1)(q-1)\right)^{1}\right\}$ and so
$E(\mathop{\mathcal{E}_{c}^{-}}(G))=q(p-2)\times 0+(q-1)(p-1)+(p-1)(q-1)=2(p-1)(q-1).$
Also, $\operatorname{L-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{1},((p-1)(q-1))^{q(p-2)},(q(p-1))^{q-1}\right\}$ and
$\operatorname{Q-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{((p-1)(q-1))^{q(p-2)},((p-1)(q-2))^{q-1},(2(p-1)(q-1))^{1}\right\}$.
We have ${|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{q(p-1)^{2}(q-1)}{2}$ and so $\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}=(p-1)(q-1)$. Therefore,
$\left|0-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=(p-1)(q-1),$
$\left|(p-1)(q-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=\left|(p-1)(q-1)-(p-1)(q-1)\right|=0,$
and
$\left|q(p-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=\left|q(p-1)-(p-1)(q-1)\right|=(p-1).$
Therefore,
$LE(\mathop{\mathcal{E}_{c}^{-}}(G))=1\times(p-1)(q-1)+q(p-2)\times 0+(q-1)\times(p-1)=2(p-1)(q-1).$
Also,
$\left|(p-1)(q-2)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|(p-1)(q-2)-(p-1)(q-1)\right|$
$=(p-1),$
and
$\left|2(p-1)(q-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2(p-1)(q-1)-(p-1)(q-1)\right|$
$=(p-1)(q-1).$
and hence
$LE^{+}(\mathop{\mathcal{E}_{c}^{-}}(G))$ $=q(p-2)\times 0+(q-1)\times(p-1)+1\times(p-1)(q-1)$
$=2(p-1)(q-1).$
@@
We have $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{t}m$ and
$E(K_{2^{t}m})=2(2^{t}m-1).$
By Theorem 6.4 we get
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-E(K_{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|})$ $=2^{t+1}(m-1)-2(m2^{t}-1)$
$=-2(2^{t}-1)<0.$

### v4

@@ 5. Realization of $\mathop{\mathcal{E}_{c}^{-}}(G)$
From this point on, we turn our attention to properties of co-Engel graphs of some non-Engel (that is, non-nilpotent groups). Our focus will be on particular examples. We remind the reader that, from this point on, we delete the isolated vertices (the elements of the Fitting subgroup).
In this section we realize certain graphs as $\mathop{\mathcal{E}_{c}^{-}}(G)$ for certain finite groups. Let $K_{n}$ be the complete graph on $n$ vertices and $K_{n_{1},n_{2},\dots,n_{m}}$ be the complete $m$-partite graph with parts of size $n_{1},n_{2},\dots,n_{m}$. We write $K_{m\cdot n}=K_{n_{1},n_{2},\dots,n_{m}}$ if $n_{1}=n_{2}=\cdots=n_{m}=n$.
We denote the dihedral group of order $2n$ by $D_{2n}$ (in the literature this group is sometimes called $D_{n}$).
@@
The set of non-isolated vertices is $G\setminus F(G)=\{x,xy,\dots,xy^{2^{t}m-1}\}$, where $F(G)$ is the Fitting subgroup. We have
@@
We also have $a^{i}b^{t}a^{j}b^{s}=a^{i+j}b^{s+tr^{j}}$. Therefore, if $i\neq 0$ then $(a^{i}b^{j})^{p}=1$. Note that the group $F_{p,q}$ is the union of its subgroups $H_{b}=\langle b\rangle=F(F_{p,q})=L(F_{p,q})$ and
@@ 6. Genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$
The genus $g(\Gamma)$ of a graph $\Gamma$ is the smallest genus of a surface in which it can be drawn without crossing edges. In this section, we determine the genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$ for the groups considered in Section 5. It is well-known that
$g(K_{n})=\left\lceil\frac{(n-3)(n-4)}{12}\right\rceil$ (1)
@@
$g(K_{m,n})=\left\lceil\frac{(m-2)(n-2)}{4}\right\rceil.$ (2)
@@
Using Theorem 5.1–Theorem 5.3 and Theorem 6.1, we get the following results.
@@
Since $\omega(\mathop{\mathcal{E}_{c}^{-}}(G))\leq 4$, by [1, Proposition 1.4] and [2, Theorem 1.2], we have $\frac{G}{Z_{\infty}(G)}\cong D_{6}$ or $A_{4}$.
Suppose $\frac{G}{Z_{\infty}(G)}\cong D_{6}$. Let $\bar{G}=\frac{G}{Z_{\infty}(G)}$ and $\bar{x}=xZ_{\infty}(G)$ for every $x\in G$. Since $\bar{G}\cong D_{6}$, we have $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\cong\mathop{\mathcal{E}_{c}^{-}}(D_{6})$. Let $\phi:\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\longrightarrow\mathop{\mathcal{E}_{c}^{-}}(D_{6})$ be a graph isomorphism. Let $a,b\in D_{6}$ such that $D_{6}=\langle a,b\rangle$ and $a^{3}=b^{2}=1$. Then $\{b,ab,a^{2}b\}$ forms a clique in $\mathop{\mathcal{E}_{c}^{-}}(D_{6})$. Therefore $\{\bar{x}_{1}=\phi^{-1}(b),\bar{x}_{2}=\phi^{-1}(ab),\bar{x}_{3}=\phi^{-1}(a^{2}b)\}$ forms a clique of $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})$. By Theorem 3.3, every element of $x_{i}Z_{\infty}(G)$ is adjacent to every element of $x_{j}Z_{\infty}(G)$, for $i,j\in\{0,1,2\},i\neq j$. Thus $K_{m,m,m}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$, where $m=|Z_{\infty}(G)|$. Therefore, by (3), we have
@@
If $m=2$, then $G$ is a non-nilpotent group of order $12$ with $|Z_{\infty}(G)|=2$. Thus $G\cong D_{12},Q_{12}$. Therefore, $\mathop{\mathcal{E}_{c}^{-}}(G)$ is planar.
If $m=3$, then $G$ is a non-nilpotent graph of order $18$ with $|Z_{\infty}(G)|=3$. Thus $G\cong C_{3}\times D_{6}$. Here $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})\cong K_{3,3,3}$ and $\gamma(K_{3,3,3})=\frac{(3-1)(3-2)}{2}=1$. Thus $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})$ is toroidal. An embedding of $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})$ on a torus is shown in Figure 1, where $C_{3}=\langle a\rangle$.
@@
Suppose $\frac{G}{Z_{\infty}(G)}\cong A_{4}$. Suppose that $|Z_{\infty}(G)|\geq 2$. Let $\bar{G}=\frac{G}{Z_{\infty}(G)}$ and $\bar{x}=xZ_{\infty}(G)$ for every $x\in G$. Since $\bar{G}\cong A_{4}$, we have $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\cong\mathop{\mathcal{E}_{c}^{-}}(A_{4})$. Let $\phi:\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\longrightarrow\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ be a graph isomorphism. As we see in Figure 2, $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ has a subgraph isomorphic to $K_{4,4}$ with parts $H=\{(2,3,4),(1,2,4),(2,4,3),(1,4,2)\}$ and $K=\{(1,2,3),(1,3,4),(1,3,2),(1,4,3)\}$. Therefore, $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})$ has a subgraph isomorphic to $K_{4,4}$ with parts
@@
Let $z\in Z_{\infty}(G),z\neq 1$. Then, by Theorem 3.3, every element of $\bar{H}\cup z\bar{H}$ is adjacent to every element of $\bar{K}\cup z\bar{K}$, showing that $K_{8,8}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$.
Thus $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq\gamma(K_{8,8})=9$, a contradiction. Thus $|Z_{\infty}(G)|=1$ and so $G\cong A_{4}$. An embedding of the graph $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ on a torus is shown in Figure 3. This completes the proof. ∎
@@
By [1, Proposition 1.4] and [2, Theorem 1.2], we have $G/Z_{\infty}(G)\cong D_{6}$ or $A_{4}$. Suppose $G/Z_{\infty}(G)\cong A_{4}$. As seen in Figure 2, $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ has a subgraph isomorphic to $K_{4,4}$. Thus $\mathop{\mathcal{E}_{c}^{-}}(G/Z_{\infty}(G))$ has a subgraph isomorphic to $K_{4,4}$. Note that $|v(\mathop{\mathcal{E}_{c}^{-}}(G/Z_{\infty}(G)))|=|v(\mathop{\mathcal{E}_{c}^{-}}(A_{4}))|=8$. Let $v(\mathop{\mathcal{E}_{c}^{-}}(G/Z_{\infty}(G)))=\{\bar{a_{1}},\bar{a_{2}},\dots,\bar{a_{8}}\}$. Then the induced subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$ by the set $\{a_{1},a_{2},\dots,a_{8}\}$ has a subgraph isomorphic to $K_{4,4}$. Thus $\bar{\gamma}(\mathop{\mathcal{E}_{c}^{-}}(G))\geq\bar{\gamma}(K_{4,4})=2$, a contradiction. Thus $G/Z_{\infty}(G)\cong D_{6}$.
Let $m=|Z_{\infty}(G)|$. As seen in the proof of Theorem 6.12, $K_{m,m,m}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$. Thus if $m\geq 3$, then $K_{6,3}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$. But $\bar{\gamma}(K_{6,3})=2$. Therefore, $m\leq 2$. If $m=1$, then $G\cong D_{6}$ and if $m=2$, then $G\cong D_{12},Q_{12}$. Note that $\mathop{\mathcal{E}_{c}^{-}}(D_{6})\cong K_{3}$ is projective. Also, from Theorem 5.1, $\mathop{\mathcal{E}_{c}^{-}}(D_{12})\cong\mathop{\mathcal{E}_{c}^{-}}(Q_{12})\cong K_{3\cdot 2}$ which is projective. This completes the proof. ∎
@@ 7. Various spectra and energies of $\mathop{\mathcal{E}_{c}^{-}}(G)$
If $m\geq 3$ is odd and $G\cong D_{2m}$ then, by Theorem 6.6, we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m}$. Therefore, using Lemma 7.1, we get the following result.
@@
Using Theorem 5.1, Theorem 5.2 and Lemma 7.2, we get the following results.
@@
We conclude this section with the following two corollaries.
@@
We have $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{t}m$ and $E(K_{2^{t}m})=2(2^{t}m-1).$ By Theorem 7.4 we get
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-E(K_{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|})=2^{t+1}(m-1)-2(m2^{t}-1)=-2(2^{t}-1)<0.$